The rectilinear motion of a particle is defined by the displacement-time
equation as x=x0+v0t+(1/2)at2
. Find the displacement and velocity at
time t=2s while x0=250mm, v0=125mm/s and a=0.5mm/s2
1
Expert's answer
2018-02-27T10:41:08-0500
Displacement=x
X= 250+(125×2)mm+(0.5×4)/2
X=501mm
Velocity = dX/dt
V= v0+2at
V= 125+2*0.5*2
V=127mm/s